The risk-free rate is 3.02% and the market risk premium is 8.92%. A stock with a β of 0.96 just paid a dividend of $2.23. The dividend is expected to grow at 24.68% for three years and then grow at 3.43% forever. What is the value of the stock?
Answer #1

Let us first compute "r" using the CAPM model.

r = risk free rate + (beta * market risk premium)

= 3.02% + (0.96*8.92%)

= 11.58%

D0 = $2.23. Now future dividends are shown below considering the growth rates:

(in $) g (in %)
D0      2.2300
D1      2.7804         24.68
D2      3.4666         24.68
D3      4.3221         24.68
D4      4.4704           3.43

Thus P4 = D4*(1+g)/(r-g)

= 4.4704*1.0343/(11.58% - 3.43%)

= $56.71

Thus price or value of stock = D1/(1+r) + D2/(1+r)^2 + D3/(1+r)^3 + D4/(1+r)^4 + P4/(1+r)^4

= 2.7804/1.1158 + 3.4666/1.1158^2 + 4.3221/1.1158^3 + 4.4704/1.1158^4 + 56.71/1.1158^4

= $47.85

Thus value of the stock = $47.85